# Reviving the Search for Indium-Based Superconductors: First-Principles Prediction of Semimetallic Superconductivity in Cubic Nd3In

### v3

@@ Abstract
The quest for materials that simultaneously exhibit superconductivity and non trivial topology has drawn significant attention in recent years, driven by their potential to host exotic quantum states. Their unique coexistence often leads to rich physics and potential applications in quantum technologies. Here, we predict cubic Nd3In as an exceptional candidate in this class, combining strong-coupling superconductivity with distinctive topological features. Using first-principles calculations, we find that the strong-coupling superconductivity in Nd3In arises primarily due to pronounced Fermi surface nesting, leading to an electron-phonon coupling constant of . Our fully anisotropic Migdal–Eliashberg analysis predicts a superconducting transition temperature
at ambient pressure, which is the highest value reported so far among cubic semimetallic superconductors. When subjected to a pressure of 15 GPa, increases further to 18 K. Beyond superconductivity, Nd3In is found to be a Weyl semimetal, as evidenced by the presence of Fermi arcs and nontrivial topological invariants, confirming its topological nature. The combination of strong-coupling superconductivity and nontrivial topological states makes Nd3In a promising candidate for quantum transport and topological quantum computation.
@@ II Methodology and Crystal structure
First-principles calculations based on density functional theory (DFT) are performed using the plane-wave pseudopotential code Quantum ESPRESSO [78]. The exchange-correlation potential is treated within the generalized gradient approximation (GGA) using the Perdew-Burke-Ernzerhof (PBE) functional. Scalar-relativistic ultrasoft pseudopotentials are used for both Nd and In atoms. The valence electron configurations considered are , , and for In, and , , , and for Nd. A kinetic energy cutoff of 75 Ry is used for the plane-wave basis set, with a charge density cutoff of 350 Ry. The Brillouin zone is sampled using a -centered k-mesh. Gaussian smearing with a width of 0.0125 Ry is applied for electronic occupations. Structural relaxations are carried out until the total energy and atomic forces converge below Ry and Ry/Bohr, respectively. Phonon dispersions, electron-phonon interactions, and superconducting parameters are evaluated using density functional perturbation theory (DFPT). For accurate estimation of superconducting properties, the EPW package is employed, which uses Wannier interpolation of electron-phonon matrix elements to efficiently calculate key quantities on dense grids [79]. The superconducting transition temperature is estimated using the Allen-Dynes modification of the McMillan formula [80]:
where denotes the logarithmic average of phonon frequencies, and is the retarded Coulomb pseudopotential. The correction factors and are expressed as
with being the mean-square phonon frequency defined by
The total electron-phonon coupling (EPC) constant is given by
where is the Eliashberg spectral function, defined as
Here, are the electron-phonon matrix elements, and are the phonon frequencies. To further investigate the electronic structure, the Fermi surface nesting function is computed as
where is the Brillouin zone volume and is the Fermi energy. Topological properties, including surface states and invariants, are computed using WANNIER90 [81] and WANNIERTOOLS [82]. Maximally localized Wannier functions are constructed with spin-orbit coupling (SOC), and surface Green’s function methods are used to resolve the surface band structure. Topological invariants are evaluated from the Wannier-based tight-binding model.
@@ III.1 Electronic properties
When SOC is included, all these band crossings are lifted, resulting in a continuous band gap between bands 48 and 49 across the Brillouin zone, as shown in Figure 4(c). Additionally, the SOC band structure exhibits double degeneracy throughout the Brillouin zone, indicative of Kramers degeneracy [89]. This degeneracy indicates the preservation of time-reversal symmetry (TRS) in the system and suggests the possibility of nontrivial topological properties [90]. These observations indicate the presence of nontrivial topological states in Nd3In, the detailed analysis of which is presented in Section D.
@@ III.2 Phonon dispersion and electron-phonon interaction
The phonon dispersion and projected phonon density of states (PHDOS) for cubic Nd3In are shown in Figure 6. The absence of imaginary frequencies throughout the Brillouin zone confirms the dynamical stability of the structure. The phonon spectrum of Nd3In can be divided into three distinct frequency regions. This distribution reflects the expected behavior that heavier atoms like Nd vibrate at lower frequencies than lighter atoms such as In. A closer inspection of the phonon dispersion reveals the presence of soft phonon modes, or Kohn anomalies, along specific high-symmetry directions particularly along the –X–M and R–X paths. Generally, these soft phonon modes are indicative of strong electron-phonon coupling [91, 92, 93, 94]. All three acoustic branches as shown in the left panel of Figure 6, exhibit noticeable softening along these directions, with the sixth vibrational mode showing pronounced softening along R–X1. We have also examined the pressure dependence of this phonon softening, as shown in Figure S1 of the Supplementary Materials [95]). The phonon dispersion weighted by mode-resolved electron-phonon coupling strength and the corresponding spectral functions were computed at 10 GPa and 15 GPa pressure. The softening intensifies under applied pressure, suggesting a possible enhancement in the superconducting critical temperature.
@@
From the EPW calculations, we obtained a refined electron–phonon coupling (EPC) constant of , indicating strong coupling and suggesting that Nd3In is a promising candidate for phonon-mediated superconductivity. The dominant contributions to the total EPC arise from soft phonon modes, as illustrated in the left panel of Figure 6, where the mode-resolved coupling strengths are projected onto the phonon dispersion. A pronounced enhancement of is observed along the –X–M and R–X1 directions, which coincide with regions of phonon softening.
The right panel of Figure 6 shows the Eliashberg spectral function and the cumulative EPC function . Analysis of reveals that phonons below 65 cm-1 contribute approximately 62.1% of the total EPC, while intermediate-frequency modes between 65 and 90 cm-1 account for 20.8%. The remaining contribution comes from higher-frequency modes.
@@ III.4 Semimetallic topology
As previously discussed, the SOC-induced electronic band structure of Nd3In suggests the presence of nontrivial topological behavior, likely arising from time-reversal symmetry (TRS). As shown in Figure 9, the three-dimensional band structure reveals a linear crossing near the –M path within the energy range of 0.10–0.25 eV, indicating the existence of a type-I Weyl point formed between the lowest conduction band and the highest valence band. Our primary focus remains on this particular crossing, which emerges in close proximity to the Fermi level. Additionally, several other Weyl points of both type-I and type-II appear at higher energies, as shown in Figure 2. However, these states lie far above the Fermi level and are therefore not discussed here, as they are unlikely to influence the transport properties significantly. To explore the topological characteristics of the type-I Weyl point near the Fermi level, we performed a detailed analysis combining Wannier-based tight-binding models and symmetry-based topological indicators. The initial step involves constructing a tight-binding Hamiltonian using maximally localized Wannier functions (MLWFs), derived from the first-principles calculations. The orbital analysis confirms that bands near the Fermi level are mainly derived from Nd orbitals. Our self-consistent field (SCF) calculations show that the total magnetization of the system converges to zero, confirming a nonmagnetic ground state for Nd3In. The combination of zero magnetization and preserved time-reversal symmetry (TRS) motivates the calculation of topological invariants to determine the system’s topological classification [108]. These invariants are computed between bands 48 and 49 using the evolution of Wannier charge centers (WCCs) [109, 110], as implemented in the WANNIERTOOLS package. Figure 10(a)–(f) displays the evolution of the Wannier charge centers (WCCs) across the six time-reversal invariant momentum (TRIM) planes. As summarized in Table 1, we observe an odd number of WCC crossings in the , , and planes, which correspond to nontrivial indices. In contrast, the , , and planes exhibit an even number of crossings, indicating trivial indices.
@@
To further elucidate the topological character, we compute the surface states on the (001) surface. These calculations reveal a rich spectrum of topological surface states (TSSs), as illustrated in Figures 11(a) and 11(c). The surface-projected band structure along the –– direction, shown in Figure 11(a), exhibits a prominent surface state crossing near the midpoint of –, within the energy window of 0.10–0.25 eV. This feature corresponds to the bulk type-I Weyl point initially observed along the –M path in the absence of spin–orbit coupling (SOC). As discussed earlier, the inclusion of SOC lifts all such Weyl crossings and opens a continuous band gap across the Brillouin zone. However, the corresponding surface states remain topologically nontrivial, as evidenced by a distinct linear crossing that bridges the valence and conduction bands within the gapped region. This bridge is highlighted in the magnified inset of Figure 11(a), underscoring the topological origin of the surface state. The existence of Weyl points in the bulk leads to topologically protected surface states called Fermi arcs that connect projections of Weyl points with opposite chirality [113]. As shown in Figure 11(b), the Fermi arcs clearly connect the projected Weyl points, thereby confirming the topologically nontrivial nature of the material. We have also investigated the surface states along the –– high-symmetry direction, as illustrated in Figure 11(c). Notably, around the point, Rashba splitting occurs near 0.30 eV, where the initially degenerate topological surface states (TSSs) split into two distinct branches [76]. A similar feature is evident along the –– path near –0.75 eV, highlighting the presence of Rashba-like TSSs that are of considerable interest in spintronics [114, 115].
@@
Before concluding, we briefly discuss the compositional variant In3Nd, with the corresponding results summarized in Supplementary Materials [95]. In3Nd crystallizes in the AuCu3-type structure with a lattice constant of 4.721 Å without SOC (4.684 Å with SOC). We find that this material also exhibits semimetallic superconductivity. Orbital-projected band structures and projected density of states without and with SOC are shown in Figures S5 and S6, respectively. Figure S5 reveals several type-I Weyl points that are gapped out upon inclusion of SOC, resulting in a continuous gap between the highest valence band and the lowest conduction band. Notably, a Weyl point is observed precisely at the Fermi level along the – path. The semimetallic nature of In3Nd is further supported by the presence of Fermi arcs, as illustrated in Figure S7. From the superconducting perspective, In3Nd exhibits a low transition temperature of K, with a superconducting gap of 0.11 meV and an electron-phonon coupling constant . This significant decrease in relative to Nd3In is primarily attributed to a substantial decrease in the total electronic density of states at the Fermi level, which decreases from 6.19 to 1.51 states/eV due to the Nd–In site interchange. The mode-resolved electron-phonon coupling strength, Eliashberg spectral function, and electronic density of states for In3Nd are detailed in Figure S8. Interestingly, this valency-exchanged phenomenon is also observed in the La–In system, where both La3In and In3La crystallize in the AuCu3-type structure. While La3In is a conventional superconductor [116], its compositional counterpart In3La exhibits semimetallic superconductivity [117, 118].

### v4

# Semimetallic Superconductivity in Cubic Nd3In: A First-Principles Insight into Indium-Based Compounds

@@ Abstract
The quest for materials that simultaneously exhibit superconductivity and nontrivial topology has drawn significant attention in recent years, driven by their potential to host exotic quantum states. Their unique coexistence often leads to rich physics and potential applications in quantum technologies. Here, we predict cubic Nd3In as an exceptional candidate in this class, combining strong-coupling superconductivity with distinctive topological features. Using first-principles calculations, we find that the strong-coupling superconductivity in Nd3In arises primarily due to pronounced Fermi surface nesting, leading to an electron-phonon coupling constant of . Our fully anisotropic Migdal–Eliashberg analysis predicts a superconducting transition temperature at ambient pressure, which is the highest value reported so far among cubic semimetallic superconductors. When subjected to a pressure of 15 GPa, increases further to 18 K. Beyond superconductivity, Nd3In is found to be a Weyl semimetal, as evidenced by the presence of Fermi arcs and nontrivial topological invariants, confirming its topological nature. The combination of strong-coupling superconductivity and nontrivial topological states makes Nd3In a promising candidate for quantum transport and topological quantum computation.
@@ II Methodology and Crystal structure
First-principles calculations based on density functional theory (DFT) are performed using the plane-wave pseudopotential code Quantum ESPRESSO [78]. The exchange-correlation potential is treated within the generalized gradient approximation (GGA) using the Perdew-Burke-Ernzerhof (PBE) functional. Scalar-relativistic ultrasoft pseudopotentials are used for both Nd and In atoms. For Nd, the 4f electrons are strongly localized and therefore excluded from the valence states. They are treated as frozen core states following the standard open-core approach, since conventional DFT does not provide a reliable description of localized f electrons[79, 80, 81]. The valence electron configurations considered are , , and for In, and , , and for Nd. A kinetic energy cutoff of 75 Ry is used for the plane-wave basis set, with a charge density cutoff of 350 Ry. The Brillouin zone is sampled using a -centered k-mesh. Gaussian smearing with a width of 0.0125 Ry is applied for electronic occupations. Structural relaxations are carried out until the total energy and atomic forces converge below Ry and Ry/Bohr, respectively. Phonon dispersions, electron-phonon interactions, and superconducting parameters are evaluated using density functional perturbation theory (DFPT). For accurate estimation of superconducting properties, the EPW package is employed, which uses Wannier interpolation of electron-phonon matrix elements to efficiently calculate key quantities on dense grids [82]. It should be noted that EPW cannot be applied within DFT+U calculations, which are often required for localized f systems. Furthermore, the use of hybrid functionals would improve the electronic description but is computationally expensive and beyond the scope of this study. The superconducting transition temperature is estimated using the Allen-Dynes modification of the McMillan formula [83]: where denotes the logarithmic average of phonon frequencies, and is the retarded Coulomb pseudopotential. The correction factors and are expressed as with being the mean-square phonon frequency defined by The total electron-phonon coupling (EPC) constant is given by where is the Eliashberg spectral function, defined as Here, are the electron-phonon matrix elements, and are the phonon frequencies. To further investigate the electronic structure, the Fermi surface nesting function is computed as where is the Brillouin zone volume and is the Fermi energy. Topological properties, including surface states and invariants, are computed using WANNIER90 [84] and WANNIERTOOLS [85]. Maximally localized Wannier functions are constructed with spin-orbit coupling (SOC), and surface Green’s function methods are used to resolve the surface band structure. Topological invariants are evaluated from the Wannier-based tight-binding model.
@@ III.1 Electronic properties
When SOC is included, a number of these Weyl crossings are lifted, modifying the bulk electronic spectrum. This effect can be clearly seen in Fig. 4(c). In the case without SOC, three band-touching points appear within the plotted energy range, dividing the spectrum into four distinct regions labeled I–IV. These touching points prevent the opening of a global gap and give rise to apparent partial-gap features. Upon inclusion of SOC, two of these Weyl points are annihilated due to SOC-induced splitting, leaving only a single Weyl point between regions I and II. As a result, the spectrum reorganizes into two main regions, and although partial gaps are present, no full insulating gap opens. The persistence of this Weyl point ensures that Nd3In retains its semimetallic character under SOC rather than becoming a topological insulator. This behavior is consistent with previous reports on SOC in topological semimetals, where SOC often eliminates some Weyl nodes but leaves others intact, thereby stabilizing the Weyl semimetal phase [92]. Additionally, the SOC band structure exhibits double degeneracy throughout the Brillouin zone, indicative of Kramers degeneracy [93]. This degeneracy indicates the preservation of time-reversal symmetry (TRS) in the system and suggests the possibility of nontrivial topological properties [94]. These observations indicate the presence of nontrivial topological states in Nd3In, the detailed analysis of which is presented in Section D.
@@ III.2 Phonon dispersion and electron-phonon interaction
The phonon dispersion and projected phonon density of states (PHDOS) for cubic Nd3In are shown in Figure 6(a)-(b). The absence of imaginary frequencies throughout the Brillouin zone confirms the dynamical stability of the structure. The phonon spectrum of Nd3In can be divided into three distinct frequency regions. This distribution reflects the expected behavior that heavier atoms like Nd vibrate at lower frequencies than lighter atoms such as In. A closer inspection of the phonon dispersion reveals the presence of soft phonon modes, or Kohn anomalies, along specific high-symmetry directions particularly along the –X–M and R–X paths. Generally, these soft phonon modes are indicative of strong electron-phonon coupling [95, 96, 97, 98]. All three acoustic branches as shown in Figure 6(a), exhibit noticeable softening along these directions, with the sixth vibrational mode showing pronounced softening along R–X1. We have also examined the pressure dependence of this phonon softening, as shown in Figure S1 of the Supplementary Materials [99]. The phonon dispersion weighted by mode-resolved electron-phonon coupling strength and the corresponding spectral functions were computed at 10 GPa and 15 GPa pressure. The softening intensifies under applied pressure, suggesting a possible enhancement in the superconducting critical temperature. Figure 6(c) shows the Eliashberg spectral function and the cumulative EPC function .
To understand the microscopic origin of the strong electron–phonon coupling, we analyzed the phonon eigenvectors of the most strongly coupled modes. Two representative modes with the highest EPC values () are highlighted in Figure 6(a). The phonon mode at the X point (Mode 1, cm-1) is dominated by the vibrations of two Nd atoms oscillating in opposite directions, while the remaining atoms remain stationary as illustrated in Figure 6(d). The next strongly coupled mode (Mode 2, cm-1), also located at the X point, involves one Nd atom and In atom vibrating out of phase with unequal amplitudes, while the other atoms remain immobile as shown in Figure 6(e). This specific relative motion of the atoms strongly perturbs the electronic states near the Fermi energy, accounting for its significant role in the EPC.
@@
From the EPW calculations, we obtained a refined electron–phonon coupling (EPC) constant of , indicating strong coupling and suggesting that Nd3In is a promising candidate for phonon-mediated superconductivity. The dominant contributions to the total EPC arise from soft phonon modes, as illustrated in Figure 6 (a), where the mode-resolved coupling strengths are projected onto the phonon dispersion. A pronounced enhancement of is observed along the –X–M and R–X1 directions, which coincide with regions of phonon softening.
Analysis of reveals that phonons below 65 cm-1 contribute approximately 62.1% of the total EPC, while intermediate-frequency modes between 65 and 90 cm-1 account for 20.8%. The remaining contribution comes from higher-frequency modes.
@@ III.4 Semimetallic topology
As previously discussed, the SOC-induced electronic band structure of Nd3In suggests the presence of nontrivial topological behavior, likely arising from time-reversal symmetry (TRS). As shown in Figure 9, the three-dimensional band structure reveals a linear crossing near the –M path within the energy range of 0.10–0.25 eV, indicating the existence of a type-I Weyl point formed between the lowest conduction band and the highest valence band. Our primary focus remains on this particular crossing, which emerges in close proximity to the Fermi level. Additionally, several other Weyl points of both type-I and type-II appear at higher energies, as shown in Figure 2. However, these states lie far above the Fermi level and are therefore not discussed here, as they are unlikely to influence the transport properties significantly. To explore the topological characteristics of the type-I Weyl point near the Fermi level, we performed a detailed analysis combining Wannier-based tight-binding models and symmetry-based topological indicators. The initial step involves constructing a tight-binding Hamiltonian using maximally localized Wannier functions (MLWFs), derived from the first-principles calculations. The orbital analysis confirms that bands near the Fermi level are mainly derived from Nd orbitals. Our self-consistent field (SCF) calculations show that the total magnetization of the system converges to zero, confirming a nonmagnetic ground state for Nd3In. To ensure that this phase is indeed the most stable, we explicitly considered ferromagnetic (FM), antiferromagnetic (AFM), and nonmagnetic (NM) configurations. The total energies of these configurations are summarized in Table A1 in the Supplementary Information. The combination of zero magnetization and preserved time-reversal symmetry (TRS) motivates the calculation of topological invariants to determine the system’s topological classification [112]. These invariants are computed between bands 48 and 49 using the evolution of Wannier charge centers (WCCs) [113, 114], as implemented in the WANNIERTOOLS package. Figure 10(a)–(f) displays the evolution of the Wannier charge centers (WCCs) across the six time-reversal invariant momentum (TRIM) planes. As summarized in Table 1, we observe an odd number of WCC crossings in the , , and planes, which correspond to nontrivial indices. In contrast, the , , and planes exhibit an even number of crossings, indicating trivial indices.
@@
To further elucidate the topological character, we compute the surface states on the (001) surface. These calculations reveal a rich spectrum of topological surface states (TSSs), as illustrated in Figures 11(a) and 11(c). The surface-projected band structure along the –– direction, shown in Figure 11(a), exhibits a prominent surface state crossing near the midpoint of –, within the energy window of 0.10–0.25 eV. This feature corresponds to the bulk type-I Weyl point observed along the –M path. As discussed previously, the inclusion of SOC lifts some of the Weyl nodes and modifies the bulk bandstructure. Although partial gaps appear, no full insulating gap opens across the Brillouin zone. Consequently, Nd3In retains its semimetallic character, and the corresponding surface states remain topologically nontrivial, as evidenced by a distinct linear crossing that connects the valence and conduction bands within the partially gapped region. This bridge is highlighted in the magnified inset of Figure 11(a), underscoring the topological origin of the surface state. The existence of Weyl points in the bulk leads to topologically protected surface states called Fermi arcs that connect projections of Weyl points with opposite chirality [117]. As shown in Figure 11(b), the Fermi arcs clearly connect the projected Weyl points, thereby confirming the topologically nontrivial nature of the material. We have also investigated the surface states along the –– high-symmetry direction, as illustrated in Figure 11(c). Notably, around the point, Rashba splitting occurs near 0.30 eV, where the initially degenerate topological surface states (TSSs) split into two distinct branches [76]. A similar feature is evident along the –– path near –0.75 eV, highlighting the presence of Rashba-like TSSs that are of considerable interest in spintronics [118, 119].
@@
Before concluding, we briefly discuss the compositional variant In3Nd, with the corresponding results summarized in Supplementary Materials [99]. The thermodynamic stability of both In3Nd and Nd3In was evaluated by constructing the convex hull of the Nd–In system as shown in Figure S5 of Supplementary Materials. In3Nd crystallizes in the AuCu3-type structure with a lattice constant of 4.721 Å without SOC (4.684 Å with SOC). We find that this material also exhibits semimetallic superconductivity. Orbital-projected band structures and projected density of states without and with SOC are shown in Figures S6 and S7, respectively. Figure S5 reveals several type-I Weyl points that are gapped out upon inclusion of SOC, resulting in a continuous gap between the highest valence band and the lowest conduction band. Notably, a Weyl point is observed precisely at the Fermi level along the – path. The semimetallic nature of In3Nd is further supported by the presence of Fermi arcs, as illustrated in Figure S8. From the superconducting perspective, In3Nd exhibits a low transition temperature of K, with a superconducting gap of 0.11 meV and an electron-phonon coupling constant . This significant decrease in relative to Nd3In is primarily attributed to a substantial decrease in the total electronic density of states at the Fermi level, which decreases from 6.19 to 1.51 states/eV due to the Nd–In site interchange. The mode-resolved electron-phonon coupling strength, Eliashberg spectral function, and electronic density of states for In3Nd are detailed in Figure S8. Interestingly, this valency-exchanged phenomenon is also observed in the La–In system, where both La3In and In3La crystallize in the AuCu3-type structure. While La3In is a conventional superconductor [120], its compositional counterpart In3La exhibits semimetallic superconductivity [121, 122].